OpenStudy (anonymous):

The perimeter of a rectangle is 420 cm. The length is 10 cm more than three times the width. Find the area of the rectangle

OpenStudy (anonymous):

420= 2l+2w 420=2(10+3w)+2w 420=20+6w+2w 400=8w w=50cm therefore l=160cm Also, Area=l x w so A= 160 x 50 A=8000cm2

OpenStudy (anonymous):

understood?

OpenStudy (anonymous):

yes i love you

OpenStudy (anonymous):

:)